The mean KE of a molecule at 0 C is 5.621 x 10^-14 ergs. Calculate no. of molecules in gm molecule of gas.

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi,

KE= 3/2 nRT= 5.621 *10^-14*6.023*10^23

n= 2/3 * 562.1*6.023/8.314*273

n=0.9944

no of molecules = 0.9944* 6.023*10^23